OpenStudy (anonymous):

please! check my answer. Find the sume of series. 2+4+8.....+ 128

OpenStudy (anonymous):

4/2= 2 , r=2 126/2 = 63 term Sn = (1/2) ( 2+ 128) Sn= 4.095 ? My anwser not match with key is: 254

OpenStudy (aravindg):

no

OpenStudy (angela210793):

u can't do 126/2

OpenStudy (aravindg):

ur answer wrong

OpenStudy (aravindg):

the formula u used is wrong

OpenStudy (aravindg):

iit is n/2(a+an)

OpenStudy (aravindg):

rate gud answer plzzz

OpenStudy (anonymous):

yes that the fomula I use it

OpenStudy (anonymous):

(63/2)*(2+126) = 4.095

OpenStudy (angela210793):

an=a1*r^(n-1) 128=2*2(n-1) 64=2^(n-1) 2^6=2^(n-1) 6=n-1 n=7 so Sn=a1*(r^n -1)/(r-1)=2*(2^7 -1)/(2-1)=2*127=254

OpenStudy (angela210793):

this is the way how u should solve it and is the right one

OpenStudy (aravindg):

answer is 254

OpenStudy (aravindg):

i got it